PLEASE PLEASE PLEASE!!!! Solve this one for me! I have been stuck on this one for what seems like forever. Please show me work, so I am may learn how to do these.
Find the distance between the points: (4, 1) and (-2, 5)
Thank you in advance!

Answer by venugopalramana(3286)

Find the distance between the points: (4, 1) and (-2, 5)
Thank you in advance!
DISTANCE BETWEEN 2 POINTS (X1,Y1) AND (X2,Y2) IS GIVEN BY
D=SQRT{(X2-X1)^2+(Y2-Y1)^2}=SQRT{(4+2)^2+(1-5)^2}=SQRT(36+16)=SQRT(52)=2SQRT(13)